OpenStudy (anonymous):

Two towns that are 31.5 km apart are 3 cm apart on a map. What is the scale of the map? A)1 cm=29.5 km B)1 cm=9.5 km C)2 cm=21 km D)2 cm=18 km This was one question that I was confused about...Please help...

OpenStudy (anonymous):

3 cm is 31.5 km apart and 1 cm is x km apart...proportion time... 3/31.5 = 1/x (cross multiply) (3)(x) = (31.5)(1) 3x = 31.5 (divide by 3 on both sides to get x by itself) 3x/3 = 31.5/3 x = 10.5 (1 cm = 10.5 km apart) which is not an answer....but , 2 cm x 10.5 km = 21. ANSWER : 2 cm = 21 km

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

Any other answers?

OpenStudy (anonymous):

I've already looked everywhere, and my mom can't help me and neither can my familila. So, this is the only answer. Duhhh...

OpenStudy (anonymous):

scale=Map distance/Ground distance =3cm:3150000=1:1050000=1cm:10.5km so from the choices it is the same as with 2cm:21km Ans Is C

OpenStudy (anonymous):

Thanx again..